HIGH RESOLUTION INFRARED SPECTRA OF THE ν_{9} AND 2ν_{4} BANDS IN METHANOL

\underline{X. WANG}, Y. MA, A. CHIROKOLAVA, and D. S. PERRY, Department of Chemistry, University of Akron, Akron, OH 44325-3601.

Jet-cooled high resolution infrared absorption spectra of the A_{2} asymmetric C-H stretch in methanol, the ν_{9} band, were recorded from 2945 to 2991 cm^{−1}. A large part of the bending overtone 2ν_{4} appears in the same region and was also recorded. The analysis has resulted in 38 subband assignments for the ν_{9} and 2ν_{4} reaching K′ up to 4 for ν_{9} and up to 2 for 2ν_{4}. A plot of the upper state torsional energies versus K′ shows the expected cosine patterns that result from the interaction of the torsion with K-rotation; however the torsional tunnelling splitting at K′= 0 is inverted for ν_{9} with the E levels below the A levels. The A and E K′= 0 subband origins for ν_{9} are 2966.6437(4) and 2952.040(3) cm^{−1}, respectively, and for 2ν_{4} are 2958.3586(11) cm^{−1} and 2957.565(6) cm^{−1}. The ν_{9} band origin (average of A and E) was found to be about 11 cm^{−1} below the estimate from low resolution spectra (2970.0 cm^{−1}). A number of perturbations in the 2ν_{4} band have been identified. The inverted torsional structure of the ν_{9} band supports the local mode Hamiltonian that was developed to explain the inverted torsional structure of the ν_{2} asymmetric C-H stretch and the regular torsional splitting of the ν_{3} symmetric C-H stretch. The model takes into account the difference in the local C-H frequency between the trans and gauche positions. The local mode parameters are the local C-H frequency ω = 2934.0 cm^{−1}, the local-local coupling parameter λ = −42.2 cm^{−1}, and the stretch-torsion coupling parameter μ = 12.9 cm^{−1}. This model yielded the correct A/E ordering and fit the 6 K′= 0 subband origins with a standard deviation of 0.4 cm^{−1}. Qualitative agreement with the K′-dependence of the torsional energies was obtained.